OpenStudy (anonymous):

A grocer wants to make a 10-pound mixture of peanuts and cashews that he can sell for $4.75 per pound. If peanuts cost $4.00 per pound and cashews cost $6.50 per pound, how many pounds of cashews should he use? 3 4 6 7

OpenStudy (anonymous):

is it 3

OpenStudy (anonymous):

say he uses x pounds of cashew. so, there is 10-x pounds of peanut the mixture costs 4(10-x)+6.5 x dollars given, 10 kg mixture is sold for $4.75 per pound so, total cost= $4.75*10=$47.50 so, 4(10-x)+6.5x=47.50 40-4x+6.5x=47.50 2.5 x=7.5 x=3 so, there is 3 pounds of cashew